(x²+2xy-7x)dx+(x²+2y²-3y)dy=0


1
Expert's answer
2021-10-07T13:20:03-0400
"M(x, y)=x^2+2xy-7x"

"\dfrac{\partial M}{\partial y}=2x"


"N(x, y)=x^2+2y^2-3y"

"\dfrac{\partial N}{\partial x}=2x"

"\dfrac{\partial M}{\partial y}=2x=\dfrac{\partial N}{\partial x}"

The differential equation "(x^2+2xy-7x)dx+(x^2+2y^2-3y)dy=0"  is an exact equation.


"\begin{cases}
 \dfrac{\partial u}{\partial x}=M(x, y) \\ \\
 \dfrac{\partial u}{\partial y}=N(x, y)
\end{cases}"

"\begin{cases}
 \dfrac{\partial u}{\partial x}=x^2+2xy-7x \\ \\
 \dfrac{\partial u}{\partial y}=x^2+2y^2-3y
\end{cases}"

"u(x, y)=\int(x^2+2xy-7x)dx+\varphi(y)"

"=\dfrac{1}{3}x^3+x^2y-\dfrac{7}{2}x^2+\varphi(y)"

"\dfrac{\partial u}{\partial y}=x^2+\varphi'(y)=x^2+2y^2-3y"

"\varphi'(y)=2y^2-3y"

"\varphi(y)=\int(2y^2-3y)dy=\dfrac{2}{3}y^3-\dfrac{3}{2}y^2-C"


"u=\dfrac{1}{3}x^3+x^2y-\dfrac{7}{2}x^2+\dfrac{2}{3}y^3-\dfrac{3}{2}y^2-C"

"\dfrac{1}{3}x^3+x^2y-\dfrac{7}{2}x^2+\dfrac{2}{3}y^3-\dfrac{3}{2}y^2=C"